Question 13 A equilateral triangular prism has an apex angle of 60.00 and is made out of flint glass and sits with its prism base on a table, as shown. A beam of red light (A 660nm) is directed at the left face of the prism. What beam angle ? relative to the horizontal will produce a beam that is parallel to the horizontal tabletop inside the prism? Enter answer here degrees
Answer #1

Question 13 A equilateral triangular prism has an apex angle of 60.00 and is made out of flint glass and sits with its prism

\frac {\sin i}{\sin 30^{\circ}}=n_{flint\: glass}=1.713

i=58.93^{\circ}

90^{\circ }-i+\phi = 60^{\circ}

\phi = 28.93^{\circ}

NOTE: I took the value of the refractive index to be 1.713. If the value is different then please change it and get the right answer.